OpenStudy (anonymous):

find exact value of 1/cos^2(40) - 1/cot(40)

OpenStudy (anonymous):

1/cos²40 - 1/cot²40 = 1/cos²40 - tan²40 =1/cos²40 - sin²40/cos²40 1 - sin²40 ------------ cos²40 Remember the Pythagorean identity Sin²x + Cos²x = 1 Cos²x = 1- Sin²x cos²40 ----------- = 1 cos²40

OpenStudy (anonymous):

thanks but the cot isnt squared

OpenStudy (anonymous):

Ahh.. I did't notice... sorry..

OpenStudy (anonymous):

you think you could do it without the cot squared?

OpenStudy (anonymous):

1/cos²40 - 1/cot40 = 1/cos²40 - tan40 =1/cos²40 - sin40/cos40 1 - sin40 cos40 ------------ cos²40 1 - sin40 -------- cos40 I don't know what's next...

OpenStudy (anonymous):

but to take out a cos wouldnt you have to have it on the sin and 1 this is why its confusing me

OpenStudy (anonymous):

Maybe you should stop there... Or inputting this on the calculator will give you 0.4663.

OpenStudy (anonymous):

it says no calculators

OpenStudy (anonymous):

but its either 1 or 0

OpenStudy (anonymous):

Ahh.. maybe its 1.. not sure... just a guess.

OpenStudy (anonymous):

ok if i change the cos^2 to 1-sin^2 could anything happen with that?

OpenStudy (anonymous):

 This part?

OpenStudy (anonymous):

at the beginning of the equation so its 1/1-sin^2(40) - tan(40)

OpenStudy (anonymous):

You'll come up with the same thing, I guess..

OpenStudy (anonymous):

man this is dumb alright thanks for the help i am gonna close it

OpenStudy (anonymous):

Okay..